A certain man has a mass of 80kg and a density of 955kg/m³ (excluding the air in his lungs).

A.) Calculate his volume.

B.) Find the buoyant force air exerts on him

C.) What is the ratio of the buoyant force of his weight?


1
Expert's answer
2021-04-26T17:12:09-0400

a)

"V=\frac{m}{\rho}=\frac{80}{955}=83.8~l,"

b)

"F=\rho_{air}gV=1.29\cdot 9.8\cdot 83.8\cdot 10^{-3}=1.06~N,"

c)

"n=\frac FP=\frac{\rho _{air} V}{m}=\frac{\rho_{air}}{\rho}=\frac{1.29}{955}=1.35\cdot 10^{-3}."